OpenStudy (anonymous):

Rewrite with no negative exponents \[\frac{ x^5 }{ x^-2 }\] the -2 is an exponent

OpenStudy (anonymous):

@bahrom7893 Hello, i posted my second question

OpenStudy (bahrom7893):

\[\frac{a}{b^c} = a*b^{-c}\]

OpenStudy (bahrom7893):

No exceptions. So in your case: \[\frac{x^5}{x^{-2}} = x^5 * x^{-(-2)} = x^5*x^2=x^{5+2}=x^7\]

OpenStudy (anonymous):

I dont get it

OpenStudy (bahrom7893):

I'll be right back, just going to go eat real quick, didn't have dinner yet.

OpenStudy (anonymous):

Okay ill be waiti

OpenStudy (bahrom7893):

Back.

OpenStudy (anonymous):

Okay

OpenStudy (bahrom7893):

Okay, so here's the rule, whenever you have a fraction as follows: \[fraction =\frac{numerator}{denominator}\]You can rewrite it this way:\[fraction=numerator*(denominator)^{-1}\]It's a rule, memorize that.

OpenStudy (bahrom7893):

So in your case, you have: \[\frac{x^{5}}{x^{-2}} = x^{5}*(x^{-2})^{-1}\]

OpenStudy (bahrom7893):

So far so good?

OpenStudy (anonymous):

yes just dont really get it still

OpenStudy (bahrom7893):

http://www.dummies.com/how-to/content/working-with-negative-exponents.html Read through this before we continue. I really am not sure how to explain this. Can't come up with an analogy :/

OpenStudy (anonymous):

Im sorry, ive read it and still dont understand this type of material?

OpenStudy (bahrom7893):

All I can say is read it again, paying more attention.

OpenStudy (anonymous):

@thivitaa hey this question

OpenStudy (thivitaa):

bah has already gave u the ans...i think..

OpenStudy (anonymous):

no he didnt

OpenStudy (anonymous):

he gave me the step right before the answer but i dont know how to finish it